# Computation of Greeks under rough Volterra stochastic volatility models using the Malliavin calculus approach

### v1

@@ Abstract
Using Malliavin calculus techniques we obtain formulas for computing Greeks under different rough Volterra stochastic volatility models. In particular we obtain formulas for rough versions of Stein-Stein, SABR and Bergomi models and numerically demonstrate the convergence.
@@ 1 Introduction
Sensitivity measures or Greeks are essential tools in hedging financial derivatives. Given a financial derivative or a portfolio with derivatives, Greeks measure the impact on its price due to changes in the different parameters on which it depends. In the simple case of Black and Scholes (1973) model, for plain vanilla options, closed-form formulas can be easily obtained. But once one departs from this simple case, that is, for more complex derivatives or for more complex models for the underlying securities, no closed-form formulas are available and the use of Monte Carlo simulations is needed. Due to the fact that numerical computation of Greeks involve the numerical computation of derivatives, procedures are complicated and slow.
@@
The extension that we are interested in the present paper is based on the fact that independent increments of the Brownian motion turned out to be a limitation in describing the real implied volatilities observed in financial markets. This helped to boost the popularity of fractional Brownian motion (fBm), a generalization of Brownian motion that allows correlation of increments depending on the so-called Hurst index $H\in(0,1)$. If $H=1/2$, one gets the standard Brownian motion, if $H>1/2$, increments are positively correlated and trajectories are more regular, and if $H<1/2$, increments are negatively correlated and we speak about the rough regime. Existence and uniqueness of a strong solution for stochastic differential equations driven by fractional Brownian motion was proved by Nualart and Ouknine (2002).
@@
The use of rough volatility models, that is, stochastic volatility models with the volatility dynamics described partially or completely by a fractional Brownian motion or a similar fractional process has become since ten years ago also a research field itself. The first rough volatility model was introduced in Alòs, León, and Vives (2007) in an effort to better describe the smirk of the implied volatility observed in markets near to expiry. In Bayer, Friz, and Gatheral (2016) and Gatheral, Jaisson, and Rosenbaum (2018) a detailed analysis of rough fractional stochastic volatility (RFSV) models is given. They seem to be very consistent with market option prices and realized volatility time series, and moreover, they provide superior volatility prediction results to several other models, see Bennedsen, Lunde, and Pakkanen (2017). Several approaches to the exact and approximate option pricing formulas in models where volatility is a fractional Ornstein–Uhlenbeck (fOU) or fractional Cox–Ingersoll–Ross (fCIR) process were introduced in Mishura (2019). A comprehensive survey of continuous stochastic volatility models including the fractional and rough models was written only recently by Di Nunno, Kubilius, Mishura, and Yurchenko-Tytarenko (2023).
@@
Computation of Greeks for jump-diffusion models, including local volatility models, is a well-developed topic, see for example Petrou (2008), Eddahbi, Lalaoui Ben Cherif, and Nasroallah (2015), Eberlein, Eddahbi, and Lalaoui Ben Cherif (2016). Two main methods are used: Fourier method and Malliavin calculus method.
@@
But for stochastic volatility models things are less developed. For Heston and Bates models we find formulas, using the Malliavin calculus approach in Davis and Johansson (2006) and in Mhlanga (2015). General formulas to compute Greeks for options under an SV model were obtained in El-Khatib (2009) using Malliavin calculus techniques and in Khedher (2012) using both Malliavin calculus techniques and a Fourier transform method. Greeks for SABR model have been obtained in Yamada (2017). Yilmaz (2018) calculated Greeks for SV models with both stochastic interest rate and stochastic volatility using Monte Carlo simulation. Finally, Yolcu-Okur, Sayer, Yilmaz, and Inkaya (2018) applied Malliavin calculus for a general SV models and calculated the Delta of different SV models such as Heston and Stein and Stein models.
@@
The structure of the paper is the following. In Section 2 we recall the definition of Greeks and Malliavin calculus fundamentals and introduce the Volterra processes. In Section 3 we introduce rough Volterra models and we derive the Malliavin weights for general Volterra stochastic volatility models. In Section 4 we derive the formulas for particular models, namely for the $\alpha$RFSV model, introduced for the first time by Merino, Pospíšil, Sobotka, Sottinen, and Vives (2021), for a new mixed $\alpha$RFSV model, and for the rough Stein-Stein model. Other models such as the rough Bergomi model, will be handled as a special case of the $\alpha$RFSV model. We will present also some numerical results, especially we will demonstrate numerically the convergence of selected obtained formulas. We conclude in Section 5.
@@ 2.1 Greeks
Consider a price process $S:=\{S_{t},t\geq 0\}.$ Let $\{{\cal F}_{t},t\geq 0\}$ the completed natural filtration generated by process $S.$ A financial derivative can be seen as a contingent claim with payoff $F$, where $F$ is a random variable adapted to ${\cal F}_{T}$ for a fixed expiry date $T>0.$ Assume a fixed instantaneous interest rate $r\geq 0.$ Under the no arbitrage principle, the price at $t\in[0,T]$ of the derivative $F$ is given by
$V_{t}:={\mathbb{E}}_{\mathbb{Q}}[\mathrm{e}^{-r(T-t)}F],$
@@
The price $V_{t}$ depends on the different parameters of the model that describes the underlying price $S$ like the price itself or its volatility, the parameters that describe the derivative like the maturity date $T$, the strike price $K$ in the case of options, and parameters of the market like the interest rate $r.$
Greeks or price sensitivities are used in hedging and for measuring and managing risk. They allow to predict the immediate future movements of a derivative price. For simplicity we assume in all the paper that $t=0$, $r\geq 0$ is a constant and $F=f(S_{T})$ for a certain function $f.$ The Greeks involved in the present paper are given by the following definition.
@@
Greeks Delta, Gamma, Rho, Vega and the derivative with respect the Hurst parameter in rough volatility models are defined respectively as
$\Delta$ $=\partial_{S_{0}}\mathbb{E}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (1)
$\Gamma$ $=\partial_{S_{0}}\Delta=\partial^{2}_{S_{0},S_{0}}\mathbb{E}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (2)
$\varrho$ $=\partial_{r}\mathbb{E}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (3)
$\mathcal{V}$ $=\partial_{V_{0}}\mathbb{E}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (4)
$\mathcal{H}$ $=\partial_{H}\mathbb{E}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (5)
where $V_{0}$ denotes the initial volatility, $r$ the fixed interest rate and $H$ the Hurst parameter associated to the fractional process underlying the rough Volterra process associated to the volatility.
@@ 2.2 Malliavin calculus and integration by parts formula
Let $W:=\{W_{t},t\in[0,T]\}$ a standard Brownian motion defined on a complete probability space $(\Omega,{\cal F},{\mathbb{P}})$. Recall that it can be seen as an isonormal Gaussian process defined on the Hilbert space $H:=L^{2}[0,T]$ and write $W=\{W(h),h\in H\}$ where
@@
Malliavin calculus is based on two dual operators, the Malliavin derivative $D$ and the Skorohod integral $\delta.$
Consider the set $\cal S$ of smooth random variables $F=f(W(h_{1}),\dots,W(h_{n}))$ where $f\in C^{\infty}_{b}({\mathbb{R}}^{n})$, the space of infinite differentiable functions such that they and all its partial derivatives are bounded, and $h_{1},\dots h_{n}$ are elements of $H.$
@@
Being $\cal S$ dense in $L^{2}$, it can be shown that derivatives $D^{k}$ are closed and densely defined operators from $L^{p}(\Omega)$ to $L^{p}(\Omega\times[0,T]^{k})$, for any $p,k\geq 1,$ with domain ${\mathbb{D}}^{k,p}$ defined as the completion of $\cal S$ by the seminorm
@@
The Malliavin derivative satisfies a chain rule in the sense that if
$\psi:{\mathbb{R}}^{n}\longrightarrow{\mathbb{R}}$
belongs to $C^{1}_{b}({\mathbb{R}}^{n})$ and $F=(F^{1},\dots,F^{n})$ with $F^{i}\in{\mathbb{D}}^{1,2}$ for any $i=1,\dots,n$, we have
@@
The adjoint of the Malliavin derivative operator $D$ is the so called divergence operator $\delta.$ It is an unbounded and closed operator from $L^{2}(\Omega\times[0,T])$ to $L^{2}(\Omega)$, densely defined on a domain $Dom(\delta)$, in such a way that for any $h\in Dom(\delta)$ and $F\in{\mathbb{D}}^{1,2}$ we have he duality relationship
@@
Proposition 2.2.
Let $F\in\mathbb{D}^{1,2}$ and $h\in Dom(\delta).$ Then $Fh\in Dom(\delta)$ and
@@
Assume $h_{s}\in{\mathbb{D}}^{1,2}$ for any $s\in[0,T]$ and the process $\{D_{t}h_{s},s\in[0,T]\}$ is Skorohod integrable. Furthermore, assume there is a version of the process
@@
Let $I$ be an open interval of ${\mathbb{R}}.$ Let $\{F^{\theta},\theta\in I\}$ and $\{\Phi^{\theta},\theta\in I\}$ be two families of functionals on ${\mathbb{D}}^{1,2}$, both continuously differentiable with respect to $\theta\in I$. Assume we have $h\in L^{2}(\Omega\times[0,T])$ such that $h_{t}\in{\mathbb{D}}^{1,2}$ for any $t\in[0,T]$,
@@
belongs to $Dom(\delta)$ and is continuous in $\theta.$ Then,
@@
for any function $f\in C^{1}_{b}(\mathbb{R})$ such that $f(F^{\theta})\in L^{2}(\Omega).$
@@
2.3 Rough Volterra processes
A Gaussian Volterra process is defined as a process $Y=\{Y_{t},t\geq 0\}$ such that it can be represented as

### v2

@@ Abstract
Using Malliavin calculus techniques, we obtain formulas for computing Greeks under different rough Volterra stochastic volatility models. Due to the fact that underlying prices are not always square integrable, we extend the classical integration by parts formula to integrable but not necessarily square integrable functionals. First of all, we obtain formulas for general stochastic volatility (SV) models, concretely the Greeks Delta, Gamma, Rho, Vega and we introduce the Greek with respect to the roughness parameter. Then, the particular case of rough Volterra SV models is analyzed. Finally, three examples are treated in detail: the family of alpha-RFSV models, that includes rough versions of SABR and Bergomi models, a mixed alpha-RFSV model with two different Hurst parameters representing short (roughness) and long memory, and the rough Stein-Stein model. For different models and Greeks we show a numerical convergence of our formulas in Monte Carlo simulations and depict for example a dependence of the Greeks on the roughness parameter.
Received 4 July 2025
@@ 1 Introduction
Sensitivity measures or Greeks are essential tools in hedging financial derivatives. Given a financial derivative or a portfolio with derivatives, Greeks measure the impact on its price as a result of changes in the different parameters on which it depends. In the simple case of Black and Scholes (1973) model, for plain vanilla options, closed-form formulas can easily be obtained. But once one departs from this simple case, that is, for more complex derivatives or for more complex models for the underlying securities, no closed-form formulas are available, and the use of Monte Carlo simulations is needed. Due to the fact that numerical computation of Greeks involves the numerical computation of derivatives, procedures are complicated and slow.
@@
The extension in which we are interested in this paper is based on the fact that independent increments of the Brownian motion turned out to be a limitation in describing the real implied volatilities observed in financial markets. This helped to increase the popularity of the fractional Brownian motion (fBm), a generalization of the Brownian motion that allows the correlation of increments depending on the so-called Hurst index $H\in(0,1)$. If $H=1/2$, one gets the standard Brownian motion, if $H>1/2$, the increments are positively correlated and trajectories are more regular, and if $H<1/2$, the increments are negatively correlated and trajectories are less regular. In the latter case we speak about the rough regime. Existence and uniqueness of a strong solution for stochastic differential equations driven by fractional Brownian motion was proved by Nualart and Ouknine (2002).
@@
The use of rough volatility models, that is, stochastic volatility models with the volatility dynamics described partially or completely by a fractional Brownian motion or a similar fractional process has since ten years ago also become a research field itself. The first rough volatility model was introduced in Alòs, León, and Vives (2007) in an effort to better describe the smirk of the implied volatility observed in markets near expiration. In Bayer, Friz, and Gatheral (2016) and Gatheral, Jaisson, and Rosenbaum (2018) a detailed analysis of rough fractional stochastic volatility (RFSV) models is given. They seem to be very consistent with market option prices and realized volatility time series, and moreover, they provide superior volatility prediction results to several other models, see Bennedsen, Lunde, and Pakkanen (2017). Several approaches to the exact and approximate option pricing formulas in models where volatility is a fractional Ornstein–Uhlenbeck (fOU) or fractional Cox–Ingersoll–Ross (fCIR) process were introduced in Mishura (2019). A comprehensive survey of continuous stochastic volatility models including the fractional and rough models was written only recently by Di Nunno, Kubilius, Mishura, and Yurchenko-Tytarenko (2023).
@@
Computation of Greeks for jump-diffusion models, including local volatility models, is a well-developed topic, see, for example, Petrou (2008), Eddahbi, Lalaoui Ben Cherif, and Nasroallah (2015), Eberlein, Eddahbi, and Lalaoui Ben Cherif (2016). Two main methods are used: Fourier transform method and Malliavin calculus method.
@@
But for stochastic volatility models things are less developed. For Heston and Bates models we find formulas, using the Malliavin calculus approach in Davis and Johansson (2006) and in Mhlanga (2015). General formulas to compute Greeks for options under an SV model were obtained by El-Khatib (2009) using Malliavin calculus techniques and in Khedher (2012) using both Malliavin calculus techniques and the Fourier transform method. Greeks for SABR model have been obtained in Yamada (2017). Yilmaz (2018) calculated Greeks for SV models with both stochastic interest rate and stochastic volatility using Monte Carlo simulation. Finally, Yolcu-Okur, Sayer, Yilmaz, and Inkaya (2018) applied Malliavin calculus for a general SV models and calculated the Delta of different SV models such as Heston and Stein and Stein models.
@@
The structure of the paper is the following. In Section 2 we recall the definition of Greeks and Malliavin calculus fundamentals and introduce Volterra processes. In Section 3 we introduce rough Volterra models and we derive the Malliavin weights for general Volterra stochastic volatility models. In Section 4 we derive the formulas for particular models, namely for the $\alpha$RFSV model, introduced for the first time by Merino, Pospíšil, Sobotka, Sottinen, and Vives (2021), for a new mixed $\alpha$RFSV model, and for the rough Stein-Stein model. Other models such as the rough Bergomi model, will be handled as a special case of the $\alpha$RFSV model. We will present also some numerical results, especially we will demonstrate numerically the convergence of selected obtained formulas. We conclude in Section 5.
@@ 2.1 Greeks
Consider a price process $S:=\{S_{t},t\geq 0\}$. Let $\{{\cal F}_{t},t\geq 0\}$ be the completed natural filtration generated by process $S.$ A financial derivative can be seen as a contingent claim with payoff $F$, where $F$ is a random variable adapted to ${\cal F}_{T}$ for a fixed expiry date $T>0.$ Assume a fixed instantaneous interest rate $r\geq 0.$ Under the no arbitrage principle, the price at $t\in[0,T]$ of the derivative $F$ is given by
$P_{t}:={\mathbb{E}}_{\mathbb{Q}}[\mathrm{e}^{-r(T-t)}F|{\cal F}_{t}],$
@@
The price $P_{t}$ depends on the different parameters of the model that describes the underlying price $S$ like the initial price $S_{0}$ or the initial volatility or variance $V_{0}$, the parameters that describe the derivative like the maturity date $T$ or the strike price $K$ in the case of options, and parameters of the market like the interest rate $r.$
Greeks or price sensitivities are used in hedging and for measuring and managing risk. They allow to predict the immediate future movements of a derivative price. For simplicity we assume in all the paper that $t=0$ and $F=f(S_{T})$ for a certain function $f.$ The Greeks involved in the present paper are given by the following definition.
@@
Greeks Delta, Gamma, Rho, Vega and the derivative with respect to the Hurst parameter in rough volatility models are defined respectively as
$\Delta$ $=\partial_{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (1)
$\Gamma$ $=\partial_{S_{0}}\Delta=\partial^{2}_{S_{0},S_{0}}\mathbb{E}_{\mathbb{Q}}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (2)
$\varrho$ $=\partial_{r}\mathbb{E}_{\mathbb{Q}}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (3)
$\mathcal{V}$ $=\partial_{V_{0}}\mathbb{E}_{\mathbb{Q}}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (4)
$\mathcal{H}$ $=\partial_{H}\mathbb{E}_{\mathbb{Q}}\left[\mathrm{e}^{-rT}f(S_{T})\right],$ (5)
where $S_{0}$ is the initial price of the stock, $V_{0}$ denotes the initial volatility, $r$ the fixed interest rate and $H$ the Hurst parameter associated to the fractional process underlying the rough Volterra process associated to the volatility.
@@ 2.2 Malliavin calculus and integration by parts formula
Let $W:=\{W_{t},t\in[0,T]\}$ be a standard Brownian motion defined on a complete probability space $(\Omega,{\cal F},{\mathbb{P}})$. Recall that it can be seen as an isonormal Gaussian process defined on the Hilbert space $H:=L^{2}[0,T]$ and write $W=\{W(h),h\in H\}$ where
@@
Malliavin calculus, sometimes also referred as Malliavin-Skorohod calculus, is based on two dual operators, the Malliavin derivative $D$ and the Skorohod integral $\delta$.
Consider the set $\cal S$ of smooth random variables $F=f(W(h_{1}),\dots,W(h_{n}))$ where $f\in C^{\infty}_{b}({\mathbb{R}}^{n})$, the space of infinitely differentiable functions such that they and all its partial derivatives are bounded, and $h_{1},\dots h_{n}$ are elements of $H.$
@@
Being $\cal S$ dense in $L^{1}$, it can be shown that derivatives $D^{k}$ are closed and densely defined operators from $L^{p}(\Omega)$ to $L^{p}(\Omega\times[0,T]^{k})$, for any $p,k\geq 1,$ with domain ${\mathbb{D}}^{k,p}$ defined as the completion of $\cal S$ by the seminorm
@@
We can also define, for $k\geq 1$, the space $\mathbb{D}^{k,\infty}=\bigcap_{p\geq 1}\mathbb{D}^{k,p}$.
Malliavin derivative satisfies a chain rule in the sense that if $\psi:{\mathbb{R}}^{n}\longrightarrow{\mathbb{R}}$ belongs to $C^{1}({\mathbb{R}}^{n})$ and $F=(F^{1},\dots,F^{n})$ with $F^{i}\in{\mathbb{D}}^{1,p}$ for a certain $p\geq 1$ and for any $i=1,\dots,n$, we have
@@
The adjoint of the Malliavin derivative operator $D$ is the so called divergence operator $\delta.$ It is an unbounded and closed operator from $L^{2}(\Omega\times[0,T])$ to $L^{2}(\Omega)$, densely defined on a domain $\operatorname{Dom}(\delta)$, in such a way that for any $h\in\operatorname{Dom}(\delta)$ and $F\in{\mathbb{D}}^{1,2}$ we have the duality relationship
@@
We now give the definitions of two vector spaces that are contained in Dom($\delta$).
Definition 2.2.
We define $\mathbb{L}^{1,2}$ as the space of processes $u\in L^{2}(\Omega\times[0,T])$ such that $u_{t}\in\mathbb{D}^{1,2}$ for all $t\in[0,T]$.
In a similar fashion as with $\mathbb{L}^{1,2}$, we define the following space.
Definition 2.3.
We define $\mathbb{L}^{1,\infty}$ as the space of processes $u\in L^{\infty}(\Omega\times[0,T])$ such that $u_{t}\in\mathbb{D}^{1,\infty}$ for all $t\in[0,T]$.
Remark 2.4.
One can indeed check that
$\mathbb{L}^{1,\infty}\subset\mathbb{L}^{1,2}\subset\operatorname{Dom}(\delta).$
The proof of this inclusions follow from Nualart (2006) (see Proposition 1.3.1 and page 42).
@@
Theorem 2.5.
Let $F\in\mathbb{D}^{1,2}$ and $h\in\operatorname{Dom}(\delta).$ Then $Fh\in\operatorname{Dom}(\delta)$ and
@@
Assume $h\in{\mathbb{L}}^{1,2}$ and the process $\{D_{t}h_{s},s\in[0,T]\}$ belongs to $\operatorname{Dom}(\delta)$ for any $t\in[0,T].$ Furthermore, assume there is a version of the process
@@
Let $I$ be an open interval of ${\mathbb{R}}.$ Let $\{F^{\theta},\theta\in I\}$ and $\{\Phi^{\theta},\theta\in I\}$ be two families of functionals on ${\mathbb{D}}^{1,2}$, both continuously differentiable with respect to $\theta\in I$. Assume $h\in\mathbb{L}^{1,2}$ such that
@@
belongs to $\operatorname{Dom}(\delta)$ and is continuous in $\theta$. Then,
@@
for any function $f\in C^{1}_{b}(\mathbb{R}).$
@@
This integration by parts formula, however, it is not enough for our purposes. The following proposition is an extension of the previous formula for a more general class of functionals.
Proposition 2.8.
Let $I$ be an open interval of ${\mathbb{R}}$. Let $\{F^{\theta},\theta\in I\}$ and $\{\Phi^{\theta},\theta\in I\}$ be two families of functionals on $\mathbb{D}^{1,1}$ and $\mathbb{D}^{1,2}$ respectively, both continuously differentiable with respect to $\theta\in I$. Assume as well that there exists $h\in\mathbb{L}^{1,\infty}$ such that
and
$\frac{\Phi^{\theta}\partial_{\theta}F^{\theta}h}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\,\mathrm{d}t}$ (13)
belongs to $\operatorname{Dom}(\delta)$ and is continuous in $\theta$. Then,
$\partial_{\theta}\mathbb{E}\left[\Phi^{\theta}f(F^{\theta})\right]=\mathbb{E}\left[f(F^{\theta})\cdot\delta\left(\frac{\Phi^{\theta}\partial_{\theta}F^{\theta}h}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\,\mathrm{d}t}\right)\right]+\mathbb{E}[f(F^{\theta})\partial_{\theta}\Phi^{\theta}]$ (14)
for any function $f\in\mathcal{C}^{1}_{b}(\mathbb{R}).$
Proof.
The idea of the proof is to approximate $F^{\theta}$ by a sequence of functionals in $\mathbb{D}^{1,2}$, applying the integration by parts formula (Theorem 2.7) to the approximating sequence and then conclude the result taking limits. Consider then, for each $\theta\in I$, a sequence $\{F^{\theta}_{n},n\geq 1\}\subseteq\mathbb{D}^{1,2}$ such that $||F^{\theta}_{n}-F^{\theta}||_{1,1}\to 0$ as $n\to\infty$. Notice that, by replacing $\{F^{\theta}_{n},n\geq 1\}$ by a proper subsequence, we can assume without loss of generality that $F^{\theta}_{n}$ and $D_{t}F^{\theta}_{n}$ converge to $F^{\theta}$ and $D_{t}F^{\theta}$ almost surely for every $t\in[0,T]$. Using the chain rule for the approximating sequence we have that
$\int_{0}^{T}D_{t}f(F^{\theta}_{n})h_{t}\,\mathrm{d}t=f^{\prime}(F_{n}^{\theta})\int_{0}^{T}D_{t}F^{\theta}_{n}h_{t}\,\mathrm{d}t.$ (15)
Now from the fact that $h\in\mathbb{L}^{1,\infty}$, $f^{\prime}$ is bounded and $\mathbb{E}\left[||DF^{\theta}_{n}-DF^{\theta}||_{L^{2}[0,T]}\right]\to 0$ as $n\to\infty$, equation (15) converges in $L^{1}(\Omega)$ to
$\int_{0}^{T}D_{t}f(F^{\theta})h_{t}\,\mathrm{d}t=f^{\prime}(F^{\theta})\int_{0}^{T}D_{t}F^{\theta}h_{t}\,\mathrm{d}t$ (16)
or, equivalently,
$f^{\prime}(F^{\theta})=\frac{\int_{0}^{T}D_{t}f(F^{\theta})h_{t}\,\mathrm{d}t}{\int_{0}^{T}D_{t}F^{\theta}h_{t}\,\mathrm{d}t}.$ (17)
Now the proof concludes in the same way as in Theorem 2.7. ∎
Notice that the previous result is useful provided such a process $h$ exists. In Section 4 we prove that in practical cases such an $h$ always exists and it can be chosen so that $||\mathbf{1}_{[0,T]}-h_{t}||_{1,1}$ is arbitrarily small.
2.3 Rough Gaussian Volterra processes
A Gaussian Volterra process is defined as a process $Y=\{Y_{t},t\geq 0\}$ that can be represented as